4. The wavelength associated with a golf ball weighing 200g and moving at a speed of 5 m/h is of the order : (a) 10-10m (c) 10-20m (d) 10-30m (e) 10-40m 

Grade:11

1 Answers

Arun
4 years ago
 
Wavelength=h/mv,
                     =6.6*10-34 j-s/ 200*10-3kg*5/3600
                     =3.3*10-31*36/5
                     =118.8*10-31/5
                     =23*10-31
                    =2.3*10-31 (ans)
                    =10-30 (ans)
 
Regards
Arun